# Exact Diagonalization, Matrix Product States and Conformal Perturbation Theory Study of a 3D Ising Fuzzy Sphere Model

### v1

@@ 7 Level mixing effects
see (Lao:2023zis, , Eq. (C.10)) and notebook for $A_{\varepsilon}(2,\mathbf{1})$. [^21] In this section we will denote $A(k,\rho)$ as $A_{\cal O}(k,\rho)$, to emphasize whose descendants are involved. The offdiagonal correction is given by
@@
is needed in both Cases (i) and (ii). Which perturbation ${\cal V}$ can achieve that? ${\cal V}=\varepsilon$ does not work because we are working at criticality and $g_{\varepsilon}$ is tuned to zero. ${\cal V}=\varepsilon^{\prime}$ does not work either because we are working at $V_{0}=4.75$ and $g_{\varepsilon^{\prime}}$ is very small, see Fig. (11), while the coefficients $B_{\varepsilon^{\prime}}\approx 0.1$ and $\tilde{B}_{\varepsilon^{\prime}}\approx 0.13$ are also small and suppress $\delta_{12}$ even further.
We next try ${\cal V}=\varepsilon^{\prime\prime}$, of dimension $\Delta_{\varepsilon^{\prime\prime}}=6.8956(43)$ Simmons-Duffin2017 . This gives larger coefficients $B_{\varepsilon^{\prime\prime}}\approx 2.79$, $\tilde{B}_{\varepsilon^{\prime\prime}}\approx 6.17$. Its needed OPE coefficients are: [^22] The observed disparity in the size of the OPE coefficients is present already in the free theory. Namely we have $\sigma=\phi$, $\varepsilon=\frac{1}{\sqrt{2}}\phi^{2}$, $\varepsilon^{\prime}=\frac{1}{\sqrt{4!}}\phi^{4}$, $\varepsilon^{\prime\prime}=\frac{1}{\sqrt{6!}}\phi^{6}$, and $f^{\rm free}_{\sigma\sigma\varepsilon^{\prime\prime}}=f^{\rm free}_{\varepsilon\varepsilon\varepsilon^{\prime\prime}}=0$, $f^{\rm free}_{\varepsilon\varepsilon^{\prime}\varepsilon^{\prime\prime}}=\sqrt{15}\approx 3.9$, $f^{\rm free}_{\varepsilon^{\prime}\varepsilon^{\prime}\varepsilon^{\prime\prime}}=8\sqrt{5}\approx 17.9$. See e.g. (Codello:2018nbe, , App. A).
@@
The last two values ning should be taken as order of magnitude estimate.
@@
for $g_{\varepsilon^{\prime\prime}}$ which realizes the needed level splitting (50). Recall that $\Delta_{1}<\Delta_{2}$. Furthermore, we see from Fig. 8 that the avoided crossing happens at a value $\Delta_{c}$ which is in between $\Delta_{1}$ and $\Delta_{2}$, in both Cases (i) and (ii). Thus we need $\delta_{11}>0$, $\delta_{22}<0$. Unfortunately, this is impossible in the discussed model where $\delta_{11}$ and $\delta_{22}$ have the same sign, both given by $g_{\varepsilon^{\prime\prime}}$ times a positive factor. Therefore, the described CPT model is unable to reproduce the precise shape of the numerical level repulsion curves under consideration. The best we could do is to choose $g_{\varepsilon^{\prime\prime}}$ negative, in which case the avoided level crossing would happen, with a roughly observed splitting, but near a $\Delta_{c}$ which is below both $\Delta_{1}$ and $\Delta_{2}$, and not in between them as seen in Fig. 8. Resolving this discrepancy is an open problem. One possibility is to include a further perturbing operator $\varepsilon^{\prime\prime\prime}$, of dimension $7.2535(51)$ Simmons-Duffin2017 . Before going in this direction it would be necessary to ascertain the OPE coefficients of $\varepsilon^{\prime\prime}$ and $\varepsilon^{\prime\prime\prime}$ with $\varepsilon$ and $\varepsilon^{\prime}$, improving on ning .
@@ 8 Conclusion and Outlook
Numerical studies of continuous phase transitions in finite volume remain a valuable means to probe Conformal Field Theories (CFTs). For 1+1D models, the spatial manifold $S^{1}$ naturally facilitates comparisons with CFT due to radial quantization. In higher dimensions, $S^{d-1}$ achieves the same purpose, but using this in practice used to pose a challenge in achieving rotational invariance. The fuzzy sphere regulator introduced by Ref. Zhu2023 has proven to be a decisive advance in this direction. By employing a rotationally invariant Hamiltonian describing interacting electrons on $S^{2}$, this approach allowed for precise comparisons with the 3D Ising CFT data known from the conformal bootstrap, and led to insights into other observables not yet accessible by the bootstrap, and into other universality classes (see the references in the introduction). In this work, we extended the original study of Zhu2023 with an eye towards critically evaluating the method’s efficacy and potential for systematic improvement.
Through the lens of Conformal Perturbation Theory (CPT), we demonstrated how to locate the critical point, determine the speed of light, and parametrize deviations between the fuzzy sphere model and the CFT. By analyzing the model away from the fine-tuned Haldane pseudopotential parameter $V_{0}=4.75$ reported in Zhu2023 , we illustrated the correlation between the size of irrelevant CPT couplings and the finite-size effects. Our work highlights that while finite-size corrections remain significant for some energy levels, CPT provides a robust framework to mitigate these effects. In fact, CPT provides to some extent an alternative to tuning the model. Previously this was noted in Lao:2023zis for the Transverse Field Ising Model on the icosahedron, and in Voinea:2024ryq for an anionic fuzzy sphere model. In both cases the model was not tuned, but a meaningful comparison to CFT was nevertheless possible after (and only after) taking into account CPT corrections.
@@
Acknowledgements.
AML acknowledges helpful discussions with G. Cuomo, A. Dey, Y.-C. He, J. Penedones, and R. Rattazzi. LH is thankful to Frédéric Mila for his support and discussions. SR is grateful to Joan Elias-Miró, Benoit Sirois, Bingxin Lao and Hugh Osborn for helpful discussions, and to the KITP and the EPFL for the hospitality. SR is especially grateful to Ning Su ning and to David Poland, Valentina Prilepina and Petar Tadić petar for communicating their unpublished results. LH was partially supported by the Swiss National Science Foundation Grant No. 212082. SR was supported in part by the Simons Foundation grant 733758 (Simons Bootstrap Collaboration). This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). AML and SR thank SCGP, Stony Brook for hospitality, where these results were first presented by AML at the workshop “Fuzzy Sphere Meets Bootstrap”, November 6-8, 2023 AMLHStonyBrookTalk .
@@ Acknowledgements.
The main source of information about the 3D Ising CFT spectrum is the conformal bootstrap studies Simmons-Duffin2017 which used $\sigma,\varepsilon$ as the external operators. Ref. Simmons-Duffin2017 reports many other operators appearing in the OPEs $\sigma\times\sigma$, $\varepsilon\times\varepsilon$, $\sigma\times\varepsilon$, whose dimensions and OPE coefficients were extracted via the extremal functional method El-Showk:2012vjm , see Tables 2-7 in that work. We do not repeat those data from Simmons-Duffin2017 here.
In a few cases we use conformal data from other conformal bootstrap studies, which are either more rigorous or more accurate than Simmons-Duffin2017 , or simply because Simmons-Duffin2017 did not have access to that part of conformal data. This additional data is shown in Table 1. Notably we use Ref. Chang:2024whx for the scaling dimensions of $\sigma,\varepsilon$ and their OPE coefficients among themselves, and Reehorst:2021hmp ; ning for $\Delta_{\varepsilon^{\prime}}$ and its OPE coefficients.
[…]
$\sigma$- 0 0.518148806(24)Chang:2024whx 0 0 0
$\varepsilon$ + 0 1.41262528(29)Chang:2024whx 1.05185373(11)Chang:2024whx 1.53244304(58)Chang:2024whx 2.40ning
$\varepsilon^{\prime}$ + 0 3.82951(61)Reehorst:2021hmp 0.05304(16)Reehorst:2021hmp 1.5362(12)Reehorst:2021hmp 7.68ning
$C$ + 4 5.022665(28)Simmons-Duffin2017 0.069076(43)Simmons-Duffin2017 0.24792(20)Simmons-Duffin2017 1.05ning
Table 1: $\mathbb{Z}_{2},\ell,\Delta$ and OPE coefficients of a few low-lying primary operators of the 3D Ising CFT which go beyond Simmons-Duffin2017 . (The OPE coefficients are denoted by $\lambda_{ijk}$ in Ref. Chang:2024whx ; Reehorst:2021hmp but the normalization is the same for scalar operators.) Boldface errors are rigorous, other errors are nonrigorous from the extremal functional method. Results from ning have no error bar.
We will also use OPE coefficients $\lambda_{TT\varepsilon}$ and $\lambda^{i}_{CC\varepsilon}$ extracted from the conformal bootstrap in Chang:2024whx and in Poland:2023vpn ; Poland:2023bny . This will be discussed in App. B.4 where these OPE coefficients will be translated into matrix elements giving diagonal shifts of $T$ and $C$ states.
@@ Appendix B CPT details
Here we will provide further details for Sec. 3. Some of these details have already been given in Lao:2023zis . However some of the discussion of Lao:2023zis is too complicated for our present purposes, since in Lao:2023zis rotation invariance on $S^{2}$ was broken to a discrete subgroup (icosahedral symmetry), which led to additional splitting effects. We will try to make this appendix self-sufficient and will also connect the discussion to the accompanying Mathematica notebook notebook which should be consulted for the expressions and for the computations. For concreteness and since it’s the main focus of this paper, the whole discussion will be limited to $d=3$ but the same theory can be developed for any $d$.
@@ B.1 ${\cal V}$ scalar primary, $\Phi$ descendants of a scalar primary
is a linear combination of tensors $\delta_{ab}\delta_{cd}$ where $abcd$ is a permutations of $\mu_{1}\mu_{2}\nu_{1}\nu_{2}$, with coefficients which depend on $\Delta_{\cal O}$ and $\Delta_{\cal V}$. The scalar product
@@
In general, the relative factors $A(k,\rho)$ on level $k$ are functions of $\Delta_{\cal O}$ and $\Delta_{\cal V}$. The notebook notebook contains the expressions and the computations of all factors with $k\leqslant 4$:
@@ B.1.1 Why Casimir eigenvalue?
Expressions for $A(k,\rho)$ in Lao:2023zis ; notebook (see (26) for $A(1,\mathbf{3})$) show that $\Delta_{\cal V}$ enters through the conformal Casimir eigenvalue $C_{\cal V}=\Delta_{\cal V}(\Delta_{\cal V}-3)$, and that furthermore
@@
There can be understood via an alternative way of computing $A(k,\rho)$, using the conformal algebra generators acting in the Hilbert space of states on the sphere Rychkov:2016iqz ; Simmons-Duffin:2016gjk . This method is a bit harder to implement in Mathematica, but it does have some merits. We will now discuss this briefly for the first level descendants.
@@
The idea now is to commute $K_{\nu}$ past ${\cal V}(n)$ towards the right, while $P_{\mu}$ towards the left, until the hit ${\cal O}$, at which point we use the primary state conditions Rychkov:2016iqz ; Simmons-Duffin:2016gjk
@@
So we replace $-[P_{\mu},[K_{\mu},{\cal V}(n)]]$ by $[{\rm Cas}_{2},{\cal V}(n)]$ which gives the Casimir eigenvalue $C_{\cal V}=\Delta_{\cal V}(\Delta_{\cal V}-3)$ and explains the second term in $A(1,\mathbf{3})$ given in (26). Two other terms produced when replacing $-P_{\mu}K_{\mu}$ by ${\rm Cas}_{2}$ vanish. The commutator with $M_{\mu\nu}M_{\mu\nu}$ vanishes as we are meant to integrate in $n$ and $\int_{n}{\cal V}(n)$ is rotationally invariant. The matrix element
@@ B.2 ${\cal U}^{(\ell)}$ spinful primary, $\Phi$ descendants of a scalar primary
The relevant three-point function has the form (Poland:2018epd, , Eq. (23))
@@
Relative correction factors for descendants up to level $k=3$ are computed in the Mathematica notebook notebook . This reproduces Eqs. (C.34)-(C.37) from Lao:2023zis . For reasons analogous to App. B.1.1, spin $\ell$ and dimension $\Delta_{\cal U}$ enter those results via the quadratic and quartic Casimir eigenvalues (see e.g. Hogervorst:2013kva ; Osborn-notes )
@@ B.3 ${\cal V}$ scalar primary, $\Phi$ descendant of the stress tensor
It is constrained by scale invariance, rotational invariance, symmetry and tracelessness of the stress tensor, symmetry under the exchange of 00 and $\infty$, and conservation of the stress tensor. It turns out that these constraints fix it uniquely in terms of $\Delta_{\cal V}$, up to an overall constant. Integrating the matrix element of the $n$-sphere gives the shift of the stress tensor level. The matrix elements for the descendants are obtained by acting on ket and bra states with $P$’s and $K$’s and commuting them past ${\cal V}(n)$, as described in App. B.1.1. One also needs to construct the projectors on various irreps into which the descendant levels split. These computations are carried out in notebook .

### v2

@@ 7 Level mixing effects
see [29, Eq. (C.10)] and [34] for $A_{\varepsilon}(2,\mathbf{1})$. [^25] In this section we will denote $A(k,\rho)$ as $A_{\cal O}(k,\rho)$, to emphasize whose descendants are involved. The offdiagonal correction is given by
@@
is needed in both Cases (i) and (ii). Which perturbation $\cal V$} can achieve that? ${\cal V}=\varepsilon$ does not work because we are working at criticality and $g_{\varepsilon}$ is tuned to zero. ${\cal V}=\varepsilon^{\prime}$ does not work either because we are working at $V_{0}=4.75$ and $g_{\varepsilon^{\prime}}$ is very small, see Fig. 11, while the coefficients $B_{\varepsilon^{\prime}}\approx 0.1$ and $\tilde{B}_{\varepsilon^{\prime}}\approx 0.13$ are also small and suppress $\delta_{12}$ even further.
We next try ${\cal V}=\varepsilon^{\prime\prime}$, of dimension $\Delta_{\varepsilon^{\prime\prime}}=6.8956(43)$ [13]. This gives larger coefficients $B_{\varepsilon^{\prime\prime}}\approx 2.79$, $\tilde{B}_{\varepsilon^{\prime\prime}}\approx 6.17$. Its needed OPE coefficients are: [^26] The observed disparity in the size of the OPE coefficients is present already in the free theory. Namely we have $\sigma=\phi$, $\varepsilon=\frac{1}{\sqrt{2}}\phi^{2}$, $\varepsilon^{\prime}=\frac{1}{\sqrt{4!}}\phi^{4}$, $\varepsilon^{\prime\prime}=\frac{1}{\sqrt{6!}}\phi^{6}$, and $f^{\rm free}_{\sigma\sigma\varepsilon^{\prime\prime}}=f^{\rm free}_{\varepsilon\varepsilon\varepsilon^{\prime\prime}}=0$, $f^{\rm free}_{\varepsilon\varepsilon^{\prime}\varepsilon^{\prime\prime}}=\sqrt{15}\approx 3.9$, $f^{\rm free}_{\varepsilon^{\prime}\varepsilon^{\prime}\varepsilon^{\prime\prime}}=8\sqrt{5}\approx 17.9$. See e.g. [65, App. A].
@@
The last two values [66] should be taken as order of magnitude estimate.
@@
for $g_{\varepsilon^{\prime\prime}}$ which realizes the needed level splitting (50). Recall that $\Delta_{1}<\Delta_{2}$. Furthermore, we see from Fig. 8 that the avoided crossing happens at a value $\Delta_{c}$ which is in between $\Delta_{1}$ and $\Delta_{2}$, in both Cases (i) and (ii). Thus we need $\delta_{11}>0$, $\delta_{22}<0$. Unfortunately, this is impossible in the discussed model where $\delta_{11}$ and $\delta_{22}$ have the same sign, both given by $g_{\varepsilon^{\prime\prime}}$ times a positive factor. Therefore, the described CPT model is unable to reproduce the precise shape of the numerical level repulsion curves under consideration. The best we could do is to choose $g_{\varepsilon^{\prime\prime}}$ negative, in which case the avoided level crossing would happen, with a roughly observed splitting, but near a $\Delta_{c}$ which is below both $\Delta_{1}$ and $\Delta_{2}$, and not in between them as seen in Fig. 8.
Resolving this discrepancy is an open problem which we leave to future work. One possibility is to include a further perturbing operator $\varepsilon^{\prime\prime\prime}$, of dimension $7.2535(51)$ [13]. Before going in this direction it would be necessary to ascertain the OPE coefficients of $\varepsilon^{\prime\prime}$ and $\varepsilon^{\prime\prime\prime}$ with $\varepsilon$ and $\varepsilon^{\prime}$, improving on [66]. Another possibility is to try non-scalar $\cal V$}. ${\cal V}=C$ would likely not work for the same reason as ${\cal V}=\varepsilon^{\prime}$: because $g_{C}$ is already constrained to be small. On the other hand, ${\cal V}=T^{\prime}$ is worth trying.
@@ 8 Conclusion and Outlook
Numerical studies of continuous phase transitions in finite volume remain a valuable means to probe Conformal Field Theories (CFTs). For 1+1D models, the spatial manifold $S^{1}$ naturally facilitates comparisons with CFT due to radial quantization. In higher dimensions, $S^{d-1}$ achieves the same purpose, but using this in practice used to pose a challenge in achieving rotational invariance. The fuzzy sphere regulator introduced by Ref. [1] has proven to be a decisive advance in this direction. By employing a rotationally invariant Hamiltonian describing interacting electrons on $S^{2}$, this approach allowed for precise comparisons with the 3D Ising CFT data known from the conformal bootstrap, and led to insights into other observables not yet accessible by the bootstrap, and into other universality classes (see the references in the introduction). In this work, we extended the original study of [1] with an eye towards critically evaluating the method’s efficacy and potential for systematic improvement.
Through the lens of Conformal Perturbation Theory (CPT), we demonstrated how to locate the critical point, determine the speed of light, and parametrize deviations between the fuzzy sphere model and the CFT. By analyzing the model away from the fine-tuned Haldane pseudopotential parameter $V_{0}=4.75$ reported in [1], we illustrated the correlation between the size of irrelevant CPT couplings and the finite-size effects. Our work highlights that while finite-size corrections remain significant for some energy levels, CPT provides a robust framework to mitigate these effects. In fact, CPT provides to some extent an alternative to tuning the model. Previously this was noted in [29] for the Transverse Field Ising Model on the icosahedron, and in [31] for an anionic fuzzy sphere model. In both cases the model was not tuned, but a meaningful comparison to CFT was nevertheless possible after (and only after) taking into account CPT corrections.
@@
Acknowledgements
AML acknowledges helpful discussions with G. Cuomo, A. Dey, Y.-C. He, J. Penedones, and R. Rattazzi. LH is thankful to Frédéric Mila for his support and discussions. SR is grateful to Joan Elias-Miró, Benoit Sirois, Bingxin Lao and Hugh Osborn for helpful discussions, and to the KITP and the EPFL for the hospitality. SR is especially grateful to Ning Su [66] and to David Poland, Valentina Prilepina and Petar Tadić [67] for communicating their unpublished results. AML and SR thank SCGP, Stony Brook for hospitality, where these results were first presented by AML at the workshop “Fuzzy Sphere Meets Bootstrap”, November 6-8, 2023 [30].
Funding information
LH was partially supported by the Swiss National Science Foundation Grant No. 212082. SR was supported in part by the Simons Foundation grant 733758 (Simons Bootstrap Collaboration). This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ Acknowledgements
The main source of information about the 3D Ising CFT spectrum is the conformal bootstrap studies [13] which used $\sigma,\varepsilon$ as the external operators. Ref. [13] reports many other operators appearing in the OPEs $\sigma\times\sigma$, $\varepsilon\times\varepsilon$, $\sigma\times\varepsilon$, whose dimensions and OPE coefficients were extracted via the extremal functional method [68], see Tables 2-7 in that work. We do not repeat those data from [13] here.
In a few cases we use conformal data from other conformal bootstrap studies, which are either more rigorous or more accurate than [13], or simply because [13] did not have access to that part of conformal data. This additional data is shown in Table 1. Notably we use Ref. [45] for the scaling dimensions of $\sigma,\varepsilon$ and their OPE coefficients among themselves, and [44, 66] for $\Delta_{\varepsilon^{\prime}}$ and its OPE coefficients.
$\cal O$} $\mathbb{Z}_{2}$ $\ell$ $\Delta$ $f_{\sigma\sigma{\cal O}}$ $f_{\varepsilon\varepsilon{\cal O}}$ $f_{\varepsilon^{\prime}\varepsilon^{\prime}{\cal O}}$
$\sigma$ - 0 0.518148806(24)[45] 0 0 0
$\varepsilon$ + 0 1.41262528(29)[45] 1.05185373(11)[45] 1.53244304(58)[45] 2.40[66]
$\varepsilon^{\prime}$ + 0 3.82951(61)[44] 0.05304(16)[44] 1.5362(12)[44] 7.68[66]
$C$ + 4 5.022665(28)[13] 0.069076(43)[13] 0.24792(20)[13] 1.05[66]
Table 1: $\mathbb{Z}_{2},\ell,\Delta$ and OPE coefficients of a few low-lying primary operators of the 3D Ising CFT which go beyond [13]. (The OPE coefficients are denoted by $\lambda_{ijk}$ in Ref. [45, 44] but the normalization is the same for scalar operators.) Boldface errors are rigorous, other errors are nonrigorous from the extremal functional method. Results from [66] have no error bar.
We will also use OPE coefficients $\lambda_{TT\varepsilon}$ and $\lambda^{i}_{CC\varepsilon}$ extracted from the conformal bootstrap in [45] and in [69, 70]. This will be discussed in App. B.4 where these OPE coefficients will be translated into matrix elements giving diagonal shifts of $T$ and $C$ states.
@@ Appendix B CPT details
Here we will provide further details for Sec. 3. Some of these details have already been given in [29]. However some of the discussion of [29] is too complicated for our present purposes, since in [29] rotation invariance on $S^{2}$ was broken to a discrete subgroup (icosahedral symmetry), which led to additional splitting effects. We will try to make this appendix self-sufficient and will also connect the discussion to the accompanying Mathematica notebook [34] which should be consulted for the expressions and for the computations. For concreteness and since it’s the main focus of this paper, the whole discussion will be limited to $d=3$ but the same theory can be developed for any $d$.
@@ B.1 $\cal V$} scalar primary, $\Phi$ descendants of a scalar primary
is a linear combination of tensors $\delta_{ab}\delta_{cd}$ where abcdabcd is a permutations of $\mu_{1}\mu_{2}\nu_{1}\nu_{2}$, with coefficients which depend on $\Delta_{\cal O}$ and $\Delta_{\cal V}$. The scalar product
@@
[…]
@@ B.1.1 Why Casimir eigenvalue?
Expressions for $A(k,\rho)$ in [29, 34] (see (26) for $A(1,\mathbf{3})$) show that $\Delta_{\cal V}$ enters through the conformal Casimir eigenvalue $C_{\cal V}=\Delta_{\cal V}(\Delta_{\cal V}-3)$, and that furthermore
@@
There can be understood via an alternative way of computing $A(k,\rho)$, using the conformal algebra generators acting in the Hilbert space of states on the sphere [46, 47]. This method is a bit harder to implement in Mathematica, but it does have some merits. We will now discuss this briefly for the first level descendants.
@@
The idea now is to commute $K_{\nu}$ past ${\cal V}(n)$ towards the right, while $P_{\mu}$ towards the left, until the hit $\cal O$}, at which point we use the primary state conditions [46, 47]
@@
So we replace $P_{\mu},[K_{\mu},{\cal V}(n)]]$ by ${\rm Cas}_{2},{\cal V}(n)]$ which gives the Casimir eigenvalue $C_{\cal V}=\Delta_{\cal V}(\Delta_{\cal V}-3)$ and explains the second term in $A(1,\mathbf{3})$ given in (26). Two other terms produced when replacing $P_{\mu}K_{\mu}$ by ${\rm Cas}_{2}$ vanish. The commutator with $M_{\mu\nu}M_{\mu\nu}$ vanishes as we are meant to integrate in $n$ and $\int_{n}{\cal V}(n)$ is rotationally invariant. The matrix element
@@ B.2 ${\cal U}^{(\ell)}$ spinful primary, $\Phi$ descendants of a scalar primary
The relevant three-point function has the form [37, Eq. (23)]
@@
Relative correction factors for descendants up to level $k=3$ are computed in the Mathematica notebook [34]. This reproduces Eqs. (C.34)-(C.37) from [29]. For reasons analogous to App. B.1.1, spin $\ell$ and dimension $\Delta_{\cal U}$ enter those results via the quadratic and quartic Casimir eigenvalues (see e.g. [71, 72])
@@ B.3 $\cal V$} scalar primary, $\Phi$ descendant of the stress tensor
It is constrained by scale invariance, rotational invariance, symmetry and tracelessness of the stress tensor, symmetry under the exchange of 0 and $\infty$, and conservation of the stress tensor. It turns out that these constraints fix it uniquely in terms of $\Delta_{\cal V}$, up to an overall constant. Integrating the matrix element of the $n$-sphere gives the shift of the stress tensor level. The matrix elements for the descendants are obtained by acting on ket and bra states with $P$’s and $K$’s and commuting them past ${\cal V}(n)$, as described in App. B.1.1. One also needs to construct the projectors on various irreps into which the descendant levels split. These computations are carried out in [34].